A cantilever is 3 meters long. The free end is on the left. On the left is a point load of 12kN. One meter from the left a bending moment of 10 kNm anticlockwise is applied.Take E = 200 GPa and I = 37 x 10-6m4

Use Macaulay' s method and take x = 0 on the left.


1
Expert's answer
2022-03-02T00:45:01-0500

Macaulay's method:

Moment equation; M(x)=-12000(x-0)1-10000(x-1)0

"EI\frac{d^2 y}{dx^2}=M(x)=-[12(x)^1+10(x-1)^0]*10^3 Nm"

integrating;

"EI\frac{dy}{dx}=-[6(x)^2+10(x-1)^1]*10^3 +A"

at x=3, "\frac{dy}{dx}=0"

"0=-[6(3)^2+10(3-1)^1]*10^3 +A"

A=74000

So,

"EI\frac{dy}{dx}=-[6(x)^2+10(x-1)-74]*10^3"

on integrating again,

"EIy=-[2(x)^3+5(x-1)^2-74x]*10^3 +B"

at x=3, y=0

"0=-[2(3)^3+5(3-1)^2-74*3]*10^3 +B"

B=-148000

So,

"EIy=-[2(x)^3+5(x-1)^2-74x+148]*10^3"

A=74000

B=-148000

Slope at free end;

"EI\frac{dy}{dx}" at x=0 is 74000

"\frac{dy}{dx}" at x=0 is "\frac{74000}{2*10^{11}*37*10^{-6}}"

=0.01 rad